The superior jumpdrive company sells jump drives for $10 each. Manufacturing cost is $2.60 per jump drive; marketing costs are $2.40 per jump; and royalty payments are 20% of the selling price. The fixed cost of preparing the jump drive is $18000. Capacity is 15000 jump drives.
Compute the breakeven point graphically & by using the conrtibution margin approach

Answer by Theo(13342)

f = fixed cost.
s = selling price per unit.
c = cost per unit.
p = profit
x = number of units.

the contribution margin approach says that:

(s - c) * x - f = p

this formula says that the selling price per unit minus the cost per unit, times the number of units, minus the fixed cost equals profit.

when you break even, profit is equal to 0, so p = 0, and the formula becomes:

(s - c) * x - f = 0

if you add f to both sides of the equation, then the equation becomes:

(s - c) * x = f

if you divide both sides of the equation by (s - c), then the equation becomes:

x = f / (s - c)

f is equal to the fixed cost of 18,000.

s is equal to the selling price of 10 dollars per unit.

c is equal to the variable cost per unit.

those costs are:

manufacturing cost of 2.60 per unit.
marketing cost of 2.40 per unit.
royalty payments of 20% of the selling price per unit.
that would be equal to 20% * 10 dollars per unit = 2 dollars per unit.

total variable costs per unit are therefore 2.60 + 2.40 + 2 = 7 dollars per unit.

at breakeven, the formula of x = f / (s - c) becomes:

x = 18000 / (10 - 7) = 18000 / 3 = 6000.

breakeven point is when the number of units sold is 6000.

the total revenue is 6000 * 10 = 60,000
the fixed cost is 18,000
the variable cost is 6000 * 7 = 42,000

the profit is equal to the total revenue minus the total cost.

total revenue is 60,000
total cost is 18,000 + 42,000 = 60,000
profit is total revenue minus total cost = 0.

the graphical solution would look like this.

$$$